OpenStudy (anonymous):

Solve for x: 7x + 2y = 5

OpenStudy (anonymous):

5-2y=7x x=(5-2y)/7

OpenStudy (anonymous):

my book only gives three exmples and none of them show how to work this problem.

OpenStudy (anonymous):

\[7x=5-2y ... x=(5-2y)/7\]

OpenStudy (anonymous):

thanks yalls!